OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of three x plus five, over seven.? f−1(x) = The quantity of seven x minus five, over three f−1(x) = The quantity of seven x plus five, over three −f−1(x) = Seven over the quantity of three x minus five f−1(x) = Seven over the quantity of five minus three x

OpenStudy (anonymous):

\[f(x)=\frac{ 3x + 5 }{ 7 }\]

OpenStudy (anonymous):

is what it looks like, correct?

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

hold up, that's not the answer

OpenStudy (anonymous):

i know i said yes lol

OpenStudy (anonymous):

answer is

OpenStudy (anonymous):

\[f^{-1}(x)=\frac{ 7x - 5 }{ 3 }\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

np, i remember this being a pain a while back

OpenStudy (anonymous):

lol can you help with 1 more??

OpenStudy (anonymous):

sure, shoot

OpenStudy (anonymous):

Given the function f(x) = The quantity of 5x minus 3, divided by 4, which of the below expressions is correct? f−1(x) = The quantity of 3 minus 5x, divided by 4 f−1(x) = The quantity of 4x plus 3, divided by 5 f−1(x) = The quantity of 4x minus 3, divided by 5. f−1(x) = The quantity of 4x minus 3, divided by 5.

OpenStudy (anonymous):

\[f(x)=\frac{ 5x - 3 }{ 4 }\]

OpenStudy (anonymous):

is the original in this case, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the answer is

OpenStudy (anonymous):

\[f ^{-1}(x)=\frac{ 5x - 3 }{ 4 }\]

OpenStudy (anonymous):

thx youre a life saver

OpenStudy (anonymous):

mistype on that last one

OpenStudy (anonymous):

is it negative 5??

OpenStudy (anonymous):

No, i messed up, it's the second one or B

OpenStudy (anonymous):

(4x+3)/5

OpenStudy (anonymous):

ok thx man

OpenStudy (anonymous):

np

OpenStudy (anonymous):

do you know the answer to this?? lol

OpenStudy (anonymous):

Calculate the average rate of change for the given function, from x = 2 to x = 7. x f(x) −2 0 2 15 7 30

OpenStudy (anonymous):

x is -2,2,7 and f(x) is 0,15,30